OpenStudy (anonymous):

find the solutions 4|2x-5|=12

OpenStudy (anonymous):

4 and 1

Parth (parthkohli):

Erm, I think you want to explain why.\[|2x - 5| = \dfrac{12}{4} = 3\]

OpenStudy (anonymous):

Then there are two possibilities: 1) 2x-5 is Positive: \[2x-5=3\] 2) 2x-5 is Negative: \[-(2x-5)=3\] Solve both of these equations for x, and you'll have your answer.